what is the answer to 4x+6=-106-10x
Answer by Alan3354(69443)

what is the answer to 4x+6=-106-10x
-------
Depends on the question.
---------------
If you want to solve for x:
4x+6=-106-10x
Add 10x to both sides
14x+6 = -106
Subtract 6
14x = -112
x = -8